OpenStudy (anonymous):

The probability that the notes level of a wide- band amplifier will exceed 2 dB is 0.05 . Find the probabilities that among 12 such amplifiers the noise level of One will exceed 2 dB

OpenStudy (anonymous):

if i help you can i get a medal

OpenStudy (kc_kennylau):

The probability of at least one amplifier exceeding 2dB is the same as 1 - the probability of no amplifier exceeding 2 dB

OpenStudy (kc_kennylau):

\[P(X\ge1)\\=1-P(X=0)\\=1-(1-0.05)^{12}\\=0.459639912...\]

OpenStudy (anonymous):

why are you soo good at math i wish i had ur math skills